OpenStudy (anonymous):

PLZ HELP PPL What is the slope of the line that passes through the points (–5, 3) and (7, 5)?

OpenStudy (aum):

\( Slope = \Large \frac{y_2-y_1}{x_2-x_1} \)

OpenStudy (anonymous):

What have you tried?

OpenStudy (anonymous):

the answers are -1, 1/6,1,6

OpenStudy (anonymous):

I don't care, what have you done?

OpenStudy (anonymous):

i came up with 2 or something

OpenStudy (anonymous):

HOW

OpenStudy (nincompoop):

I knew you wanted that Purple Honorary title

OpenStudy (anonymous):

cuz im a bad butt thats how

OpenStudy (anonymous):

Yeah, pretty much \[m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (dannyrod2000):

CONGRATS @iambatman 700 FAN!!

OpenStudy (dannyrod2000):

lol sorry to interrupt

OpenStudy (anonymous):

Lol thanks, and don't worry this guy just wants a direct answer, so w e.

OpenStudy (dannyrod2000):

i know him :P

OpenStudy (dannyrod2000):

they deleted one of his posts

OpenStudy (dannyrod2000):

hero did

OpenStudy (anonymous):

i dont want it

OpenStudy (dannyrod2000):

oh and he wanted to become A MOD lol

OpenStudy (anonymous):

All this time you spent asking for the answer in the chat, you could've been done by now.

OpenStudy (anonymous):

Nicely done, @iambatman.

OpenStudy (anonymous):

\[\huge m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\] where m is your slope and your coordinates represent \[(x_1,y_1)~~(x_2,y_2)\]